OpenStudy (anonymous):

In a right triangle, the acute angles measure x+15 and 2x degrees. What is the measure of the larger acute angle of the triangle? 15 25 40 50

OpenStudy (jadeishere):

The angles in a triangle sum to 180 degrees. So as the right angle is 90, the other two must total 180 - 90 = 90. So, x+15 + 2x = 90 Therefore 3x = 90 - 15 = 75 and so x=25. So the smaller angles are x+15 = 40 degrees and 2x=50 degrees. So the larger one is 50 degrees.